OpenStudy (chris215):

Find dy/dx if f(x) = (x + 1)^(2x)

OpenStudy (anonymous):

method one take the log take the derivative of the log after "simplifying using property of the log then multiply your result by the original function

OpenStudy (anonymous):

method two rewrite \[f(x)=(x+1)^{2x}\] as, by definition \[\large e^{2x\ln(x+1)}\] then take the derivative of that using the chain rule all the actual work will be the same

OpenStudy (anonymous):

if you do it both ways you see that the real work is finding the derivative of \[2x\ln(x+1)\] using the product rule

OpenStudy (anonymous):

you good from there?

OpenStudy (chris215):

I got (2ln(x+1)+2x/x+1)(x-1)^(2x)

OpenStudy (anonymous):

me too

OpenStudy (chris215):

thank you so much!

OpenStudy (anonymous):

yw